OpenStudy (anonymous):

The function p(d)=1+ d/33 gives the pressure, in atmospheres(atm), at a depth d in the sea (d is in feet). Note that p(0)=1atm,p(33)=2, and so on. Find the pressure at 120 feet

OpenStudy (anonymous):

4.63

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

is it not p(120) = 1 +120 /33

OpenStudy (anonymous):

p(120) = 1 + 120/33 = 153/33 = 51/11 atmospheres.